Ps. 2o4 I C Cに1ド Eは Eご15V (a ter d) Fcwol the emea
media%2F188%2F18826dc0-8add-4048-ac67-8f
Answer #1

1. (a) C1, C2 and C3 are in parallel.

C = C1 + C2 + C3 = 30 uF

now C and C4 are in series,

Ceq = C C4 / (C + C4) = 15 uF

(b) Q4 = Q = CEq E = 225 uC

Q1 = Q2 = Q3 = 225/3 = 75 uC

(C) V1 = V2 = V3 = 75uC / 10uF = 7.5 Volt

V4 = 225 / 30 = 7.5 Volt


(d) U = C V^2 /2

U1 = U2 = U3 = 281 x 10^-6 J

U4 = 843.7 x 10^-6 J